A sample of 0.892 g of potassium chloride (KCl) is dissolved in 54.6 g of water (H2O). Determine the percent mass of KCl in the solution and its molality if the molar mass of KCl is 74.5513 grams per mole.


A sample of 0.892 g of potassium chloride (KCl) is dissolved in 54.6 g of water. What is the percent by mass of KCl in the solution?


A sample of 0.892 g of potassium chloride (KCl) is dissolved in 54.6 g of water. What is the percent by mass of KCl in the solution?


The density of a 2.45 M aqueous solution of methanol (CH3OH) is 0.976 g/mL. What is the molality of the solution? The molar mass of methanol is 32.04g.


1
Expert's answer
2022-04-22T11:25:02-0400

"W(KCl)=m(KCl)\/(m(KCl)+m(H_2O))=0.892\/(0.892+54.6)=0.016"

n(KCl)=0.892/74.5513=0.012 mol

"c(KCl)=n(KCl)\/m(H_2O)=0.012\/0.0546=0.2191mol\/kg"


2.

mass of 1l of solution is 976 g=0.976 kg. It contains 2.45 mol of methanol.

Mass of methanol is 976-32.04x 2.45=897.5 g=0.897 kg of water

So c=2.45/0.897=2.73 mol/kg